OpenStudy (anonymous):

In the circuit below R1 = R2. What is the value for the resistors if the circuit dissipates 65.4 mW?

OpenStudy (anonymous):

circuit diagram?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

R1=R2, so effective resistance(R) in this parallel circuit=R/2 V=18V Power=65.4 mW=0.0654W \[power, P=\frac{V^2}{R}\] \[R=\frac{V^{2}}{P}\] \[R=\frac{18*18}{0.0654}=4954.13\] \[R1=R2=2R=4954.13*2=9908.26 \Omega \]

OpenStudy (anonymous):

http://library.thinkquest.org/10796/ch14/ch14.htm